# Ram-EHR: Retrieval Augmentation Meets Clinical Predictions on Electronic Health Records

### v1

@@ Abstract
[…]
@@ 3.3 Augmenting Patient Visits with Summarized Knowledge via Co-training
Recall that patient visits and summarized knowledge encode complementary information for clinical prediction tasks — visits capture cooccurence relationships, while summarized knowledge encodes semantic information. To effectively aggregate these two types of information, we design a co-training approach, detailed as follows.
@@
$\mathcal{L}=$ $\mathbb{E}_{(V_{i},y_{i})\sim\mathcal{P}}~{}\ell(\widehat{y}_{i,1},y_{i})+\ell(\widehat{y}_{i,2},y_{i})$ (5)
$+\lambda\left(\mathcal{D}_{\operatorname{KL}}(\widehat{y}_{i,1},\widetilde{y})+\mathcal{D}_{\operatorname{KL}}(\widehat{y}_{i,2},\widetilde{y})\right),$
@@ 4.1 Experiment Setups
We are aware that there are other advanced clinical predictive models (Luo et al., 2020; Gao et al., 2020; Lu et al., 2021), however, these works have different focuses compared with Ram-EHR, which is of separate interest. We also refrain from directly leveraging LLMs as disease predictors due to the low inference speed as well as privacy concerns of putting patient visits into LLMs [^1] https://physionet.org/about/licenses/physionet-credentialed-health-data-license-150/.
@@
Implementation Details. In this work, we use Dragon (Lin et al., 2023) as the retriever for both passage $R_{D}(\cdot)$ and queries $R_{Q}(\cdot)$ [^2] https://huggingface.co/facebook/dragon-plus-{query,context}-encoder. We set $k=5$ during the retrieval stage. We choose UMLS-BERT (Michalopoulos et al., 2021) as the augmented model $g_{\phi}$ with 110M parameters for Ram-EHR and relevant baselines as $g_{\phi}$, with a maximum length of 512, and HyGT (Cai et al., 2022) as $f_{\theta}$ in main experiments. We also evaluate the effects of $g_{\phi}$ on PubmedBERT (Tinn et al., 2021), Clin-MobileBERT, and Clin-DistilBERT (Rohanian et al., 2023), as well as the effects of $f_{\theta}$ on GCT (Choi et al., 2020) and Transformer (Li et al., 2020) in Section 4.3. We set the learning rate to 5e-5 for $g_{\phi}$ and 1e-4 for $f_{\theta}$, batch size to 32, and the number of epochs to 5. We select $\beta,\lambda$ based on the performance of the validation set, and present the parameter study in Appendix 4.3.4. All the experiments are conducted on a Linux server with one NVIDIA A100 GPU.

### v2

@@ Abstract
Ran Xu1thanks:  Equal contribution., Wenqi Shi211footnotemark: 1, Yue Yu2, Yuchen Zhuang2, Bowen Jin3 May D. Wang2, Joyce C. Ho1, Carl Yang1 1 Emory University  2 Georgia Institute of Technology 3 University of Illinois at Urbana Champaign {ran.xu,joyce.c.ho,j.carlyang}@emory.edu, {wshi83,yczhuang,yueyu,maywang}@gatech.edu, bowenj4@illinois.edu
@@ 3.3 Augmenting Patient Visits with Summarized Knowledge via Co-training
Recall that patient visits and summarized knowledge encode complementary information for clinical prediction tasks — visits capture co-occurrence relationships, while summarized knowledge encodes semantic information. To effectively aggregate these two types of information, we design a co-training approach, detailed as follows.
@@
$\mathcal{L}_{\text{aug}}=$ $\mathbb{E}_{(V_{i},y_{i})\sim\mathcal{P}}~{}\ell(\widehat{y}_{i,1},y_{i})+\lambda\mathcal{D}_{\operatorname{KL}}(\widehat{y}_{i,1},\widetilde{y}),$ (5)
$\mathcal{L}_{\text{loc}}=$ $\mathbb{E}_{(V_{i},y_{i})\sim\mathcal{P}}~{}\ell(\widehat{y}_{i,2},y_{i})+\lambda\mathcal{D}_{\operatorname{KL}}(\widehat{y}_{i,2},\widetilde{y}),$
@@ 4.1 Experiment Setups
Implementation Details. In this work, we use Dragon (Lin et al., 2023) as the retriever for both passage $R_{D}(\cdot)$ and queries $R_{Q}(\cdot)$ [^1] https://huggingface.co/facebook/dragon-plus-{query,context}-encoder. We set $k=5$ during the retrieval stage. We choose UMLS-BERT (Michalopoulos et al., 2021) as the augmented model $g_{\phi}$ with 110M parameters for Ram-EHR and relevant baselines as $g_{\phi}$, with a maximum length of 512, and HyGT (Cai et al., 2022) as $f_{\theta}$ in main experiments. We also evaluate the effects of $g_{\phi}$ on PubmedBERT (Tinn et al., 2021), Clin-MobileBERT, and Clin-DistilBERT (Rohanian et al., 2023), as well as the effects of $f_{\theta}$ on GCT (Choi et al., 2020) and Transformer (Li et al., 2020) in Section 4.3. We set the learning rate to 5e-5 for $g_{\phi}$ and 1e-4 for $f_{\theta}$, batch size to 32, and the number of epochs to 5. We select $\beta,\lambda$ based on the performance of the validation set. All the experiments are conducted on a Linux server with one NVIDIA A100 GPU.
@@ 5 Conclusion
Acknowledgement
We thank the anonymous reviewers and area chairs for valuable feedbacks. This research was partially supported by the internal funds and GPU servers provided by the Computer Science Department of Emory University. JH was supported by NSF grants IIS-1838200 and IIS-2145411.